OpenStudy (anonymous):

y^-2= x^-3 is there a special rule to get rid of the -3 so i can just have something =x?

OpenStudy (phi):

you could use this rule \[ \left( a^b\right)^c = a^{bc}\]

OpenStudy (phi):

in other words find the "c" that makes -3*c = 1 then raise both sides to that power (to keep it an equation)

OpenStudy (anonymous):

wow! that was so obvious haha thnx!